Using the number 32 and 13. At some point during the lives of you and the youngest person, your age will be three times his/her age at that moment. Write an equation which models how old in years each of you will be when you are three times as old as the younger person.

Answer by richwmiller(17219)

32+x=(13+x)*3
3.5 years ago
28.5=9.5*3